### 02 Apr 2026 v3

# PyEncode: An Open-Source Library for Structured Quantum State Preparation

Abstract
Quantum algorithms require encoding classical vectors as quantum states, a step known as amplitude encoding. General-purpose state preparation routines accept any input vector of length $N=2^{m}$ and produce circuits with $\mathcal{O}(2^{m})$ gates. However, vectors arising in scientific and engineering applications often exhibit mathematical structure that admits far more efficient encoding.
Recent theoretical work has established closed-form circuits for several structured vector classes, but without open-source implementations. We present PyEncode, an open-source Python library that implements this body of theory in a unified, immediately deployable framework. The library covers sparse, step, square (general interval), Walsh, geometric, and Fourier patterns, and supports weighted superpositions of pattern states via the linear combination of unitaries (LCU) protocol, enabling exact preparation of piecewise-structured vectors such as multi-interval Hamiltonians.
PyEncode exposes a single function encode(VectorObj, N) that maps a typed parameter declaration directly to a verified Qiskit circuit, with no vector materialization and no approximation. Sparse, step, and Walsh vectors require only $\mathcal{O}(m)$ gates; geometric (exponential-decay) vectors require $\mathcal{O}(m)$ gates with zero two-qubit gates; square (general interval) vectors require $\mathcal{O}(m^{2})$ gates via a QFT-based constant adder, with $\mathcal{O}(m)$ special cases; Fourier (sinusoidal) vectors require $\mathcal{O}(m^{2})$ gates via the inverse Quantum Fourier Transform — all exponentially fewer than the $\mathcal{O}(2^{m})$ cost of general state preparation. LCU combines $r$ component circuits whose total gate cost is the sum of individual component costs, with success probability $p\in(0,1]$ determined analytically. The library is available at https://github.com/UW-ERSL/PyEncode.
1 Motivation
Quantum linear solvers such as HHL [12] and QSVT [7, 18] promise exponential speedup for solving $Au=f$, but require the right-hand side $f$ to be loaded as a quantum state $|\psi_{f}\rangle$. This state preparation is a well-known bottleneck [1]: for a vector of length $2^{m}$, general state preparation requires $\mathcal{O}(2^{m})$ gates, erasing any downstream speedup.
The same bottleneck appears across quantum chemistry, where fault-tolerant phase estimation requires the PREP oracle to amplitude-encode the Pauli coefficient vector of the Hamiltonian [5, 3], and in quantum Monte Carlo, where amplitude estimation achieves a quadratic speedup only if the probability distribution can be loaded efficiently [19, 13].
However, what is common to all of these settings is that the vector to be encoded is rarely arbitrary. In computational mechanics, $\mathbf{f}$ is a discretized load or source term with a known mathematical form — a sinusoidal mode, a uniform pressure, a point force. In quantum chemistry, the Pauli coefficient vector of a translationally invariant Hamiltonian takes only a small number of distinct values. In quantum finance, the discretized probability distribution is piecewise constant.
Several authors have established that structured classical vectors admit exact, exponentially compact quantum circuits — but without open-source implementations. PyEncode fills this gap: a single function encode(VectorObj, N) maps a typed parameter declaration directly to a verified Qiskit circuit, with no vector materialization and no approximation. The LCU constructor further enables exact weighted superpositions of pattern states, with analytically determined success probability.
For example, consider a basis vector on $N=2^{6}=64$ nodes with a single nonzero entry at index $19$, i.e., $\mathbf{f}=\mathbf{e}_{19}$. Qiskit’s StatePreparation produces 97 gates. PyEncode recognizes that $\mathbf{f}$ is a computational-basis state and writes $19=(010011)_{2}$ directly in binary, reducing the circuit to 3 gates — a $32\times$ reduction with zero approximation error. [^1] Qiskit StatePreparation decomposed with reps=3; both circuits transpiled to $\{\textsc{cx},U\}$ at optimization_level=3 using Qiskit 2.3.1. The state preparation in PyEncode is always exact: no approximation is introduced.
Amplitude encoding — the encoding scheme targeted by PyEncode — is the most compact in terms of qubit count ($m=\log_{2}N$ qubits for a vector of length $N$) and is the natural choice when the algorithm requires only a property of the encoded vector, such as an inner product, norm, or expectation value [26]. It is used in quantum linear solvers [12], quantum Monte Carlo [19], and Pauli-operator preparation in quantum chemistry [5].
2 Prior Work
General-purpose state preparation
The problem of preparing an arbitrary $N=2^{m}$-dimensional quantum state has been studied extensively. Shende, Bullock, and Markov [23] established the $\mathcal{O}(2^{m})$ lower bound on gate count and gave an explicit constructive procedure. Araujo et al. [2] reduce circuit depth to $\mathcal{O}(\log^{2}N)$ via a divide-and-conquer strategy, at the cost of $\mathcal{O}(N)$ ancilla qubits. Gui et al. [11] propose a deterministic $\mathcal{O}(\log N)$-depth protocol with optimal spacetime allocation $\mathcal{O}(N)$. All of these methods treat the input vector as an opaque array and do not exploit any structure in the amplitudes.
Approximate state preparation
Several methods achieve sub-exponential complexity by allowing a controlled approximation error. Welch et al. [27] established the connection between Walsh functions and diagonal unitary circuits, showing that truncated Walsh–Fourier series yield approximately minimal-depth circuits. Holmes and Matsuura [14] showed that smooth, differentiable functions admit linear-depth circuits via matrix product state (MPS) approximations. O’Brien and Sünderhauf [22] achieve efficient approximate state preparation for piecewise-defined functions — whose amplitudes are well approximated by piecewise polynomials — via quantum singular value transformation (QSVT). Marin-Sanchez et al. [17] reduce the Gröver–Rudolph circuit complexity from $\mathcal{O}(2^{n})$ to $\mathcal{O}(2^{k_{0}(\epsilon)})$ for smooth, real-valued functions. Zylberman and Debbasch [29] introduce the Walsh Series Loader (WSL), achieving circuit depth $\mathcal{O}(1/\sqrt{\epsilon})$ independent of the number of qubits. All of these methods are approximate, working from either the numerical vector or a smoothness assumption. PyEncode targets the complementary regime: vectors whose structure is declared a priori, for which it produces exact, closed-form circuits with zero approximation error.
Structure-exploiting state preparation
A smaller body of work exploits specific structural properties to reduce gate complexity. Grover and Rudolph [10] showed that if the cumulative amplitudes $\sum_{i=0}^{k}f_{i}^{2}$ can be computed efficiently — as they can for any analytically defined vector — the state can be prepared with $\mathcal{O}(m)$ circuit depth via a sequence of controlled $R_{y}$ rotations determined by cumulative integrals. However, the full binary tree structure means the gate count remains $\mathcal{O}(2^{m})$, whereas PyEncode’s synthesizers exploit the algebraic structure of each pattern directly — such as the binary decomposition underlying STEP — to achieve $\mathcal{O}(\mathrm{poly}(m))$ gate count and, for several patterns, depth 1. Gleinig and Hoefler [8] give an exact $\mathcal{O}(sm)$-gate algorithm for sparse states with $s$ nonzero coefficients. Shukla and Vedula [24] derive closed-form circuits for interval uniform superpositions — quantum states encoding a uniform amplitude over a contiguous index range — and establish their $\mathcal{O}(m)$ gate complexity using multi-controlled operations and the binary structure of the interval endpoints. PyEncode’s SQUARE synthesizer extends this to general intervals $k_{1},k_{2})$ via a Draper QFT-based constant adder [6]. PyEncode’s FOURIER synthesizer prepares sinusoidal vectors exactly by loading their DFT coefficients as a SPARSE state and applying the inverse QFT [21].
PyEncode’s contribution
The works cited above provide theoretical constructions without open-source implementations. PyEncode fills this gap, providing a unified, immediately deployable library that implements these constructions directly: SPARSE via Gleinig–Hoefler [8]; STEP via Shukla–Vedula [24]; WALSH via Welch et al. [27]; and FOURIER via the exact, finite-term inverse-QFT pipeline of Gonzalez-Conde et al. and Moosa et al. [9, 20]. The SQUARE synthesizer extends STEP to general intervals $k_{1},k_{2})$ by composing a Shukla–Vedula step circuit with a Draper QFT-based constant adder [6], at $\mathcal{O}(m^{2})$ cost in general and $\mathcal{O}(m)$ for power-of-2-aligned blocks. The GEOMETRIC synthesizer packages the well-known product-state structure of $r^{i}$ [28] as a named pattern, preparing exponential-decay sequences in $\mathcal{O}(m)$ single-qubit gates with zero entangling gates. The LCU constructor implements LCU [3] for weighted superpositions of pattern states, with analytically determined success probability.
3 Structured Encoding
PyEncode exposes a single function
encode(VectorObj, N, validate=False, tol=1e-6)
where VectorObj is a typed constructor that declares the mathematical form of the vector to be encoded, and N is the vector length ($N=2^{m}$, $m$ qubits). The function returns (circuit, info); full details of the EncodingInfo dataclass fields are described in Section 3.8. No classical vector is ever materialized during synthesis. All patterns encode real-valued amplitude vectors; complex-valued state preparation (e.g., via $R_{y}$$R_{z}$ pairs in the Gleinig–Hoefler tree) is a straightforward extension but lies outside the current scope.
All examples below import:
from pyencode import encode, SPARSE, STEP, SQUARE, FOURIER, WALSH, GEOMETRIC, LCU
Table 1 summarizes the seven supported patterns.
Table 1: Recognized vector patterns in PyEncode. SQUARE uses a Draper QFT-based constant adder [6]: $\mathcal{O}(m)$ for $k_{1}=0$ or power-of-2-aligned blocks; $\mathcal{O}(m^{2})$ in general.
Pattern Constructor Form $\mathcal{O}(\cdot)$ Source
Sparse SPARSE([(x,a),...]) $\sum_{j}\alpha_{j}|x_{j}\rangle$ $\mathcal{O}(sm)$ [8]
Step STEP(k_s,c) $c\,\mathbf{1}[i<k_{s}]$ $\mathcal{O}(m)$ [24]
Square SQUARE(k1,k2,c) $c\,\mathbf{1}[k_{1}\leq i<k_{2}]$ $\mathcal{O}(m^{2})^{\dagger}$ [24, 6]
Walsh WALSH(k,c_pos,c_neg) $c_{\pm}$ on $b_{k}(i)=0/1$ halves $\mathcal{O}(m)$ [27]
Geometric GEOMETRIC(ratio) $c\,\mathrm{ratio}^{\,i}$ $\mathcal{O}(m)$ [28]
Fourier FOURIER(modes=[...]) $\sum_{t}A_{t}\sin(2\pi n_{t}i/N+\varphi_{t})$ $\mathcal{O}(m^{2})$ [21]
LCU LCU([(w,VObj),…]) $\sum_{j}w_{j}|f^{(j)}\rangle$ $\mathcal{O}(\sum_{i}C_{i})$ [3]
3.1 Sparse
A superposition of $s$ basis states at explicitly declared indices with (possibly distinct) real amplitudes, as defined by Gleinig and Hoefler [8]:
$|\psi\rangle=\frac{1}{\mathcal{N}}\sum_{j=1}^{s}\alpha_{j}\,|x_{j}\rangle,\quad S=\{x_{1},\ldots,x_{s}\}\subseteq\{0,\ldots,N{-}1\}$ (1)
where $\mathcal{N}$ is the normalization constant. Constructor: SPARSE([(x1, alpha1), (x2, alpha2), ...]). Circuit complexity: $\mathcal{O}(s\,m)$ [8]. The Gleinig–Hoefler algorithm builds the state recursively by partitioning $S$ along qubit boundaries; at each level a single controlled-$R_{y}$ rotation selects the relative weight between the two halves. The total gate count is $\mathcal{O}(s\,m)$ for arbitrary $s$ and $N=2^{m}$.
Example. The basis vector $\mathbf{e}_{19}$ on $N=64$ nodes is the $s=1$ special case of Gleinig–Hoefler [8]. PyEncode writes $19=(010011)_{2}$ and applies $X$ gates to qubits where the bit is 1:
circuit, info = encode(SPARSE([(19, 1.0)]), N=64)
# info.complexity -> "O(s*m)"
# info.gate_count -> 3 (Hamming weight of 19)
Figure 1 shows the vector and circuit.
Figure 1: Left: Sparse ($s=1$): $\alpha_{19}=1$, N=64N=64Right: PyEncode circuit.
Multi-entry example.
circuit, info = encode(SPARSE([(1, 3.0), (6, 4.0)]), N=8)
# info.complexity -> "O(s*m)" (s=2)
# info.gate_count -> 5
Figure 2 shows the two-entry sparse vector.
Figure 2: Left: Sparse ($s=2$): $\alpha_{1}=3,\;\alpha_{6}=4$, N=8N=8Right: PyEncode circuit.
3.2 Step
Constant amplitude $c$ on the prefix $0,k_{s})$, zero otherwise:
$f_{i}=c\,\mathbf{1}[i<k_{s}]$ (2)
Constructor: STEP(k_s, c). Circuit complexity: $\mathcal{O}(m)$. This is the interval uniform superposition studied by Shukla and Vedula [24]. Their construction exploits the binary representation of $k_{s}$ to decompose the multi-controlled preparation into $\mathcal{O}(m)$ elementary gates. PyEncode implements this construction directly. Arises as a one-sided load distribution in structural mechanics and as the PREP state for prefix-structured Hamiltonian coefficients. The special case $k_{s}=N$ covers the entire range, producing the uniform superposition $H^{\otimes m}|0\rangle$ in $m$ gates.
Example.
circuit, info = encode(STEP(k_s=4, c=1.0), N=8)
# info.complexity -> "O(m)"
# info.gate_count -> 2
Figure 3 shows the step vector and its $\mathcal{O}(m)$ circuit.
Figure 3: Left: Step: $f_{[:4]}=1$, N=8N=8Right: PyEncode circuit.
3.3 Square
Constant amplitude $c$ on a general interval $k_{1},k_{2})$, zero otherwise:
$f_{i}=c\,\mathbf{1}[k_{1}\leq i<k_{2}]$ (3)
Constructor: SQUARE(k1, k2, c). Arises as a uniform patch load in structural mechanics.
3.3.1 Construction
The interval $k_{1},k_{2})$ is a shift of the prefix interval $[0,w)$ by the classical constant $k_{1}$, where $w=k_{2}-k_{1}$. The circuit proceeds in two ancilla-free stages:
  1. 1.
    STEP(w)(w): prepare the uniform superposition over $[0,w)$ using the Shukla–Vedula step circuit [24] at $\mathcal{O}(m)$ cost.
  2. 2.
[…]
The result is the uniform superposition $\frac{1}{\sqrt{w}}\sum_{i=k_{1}}^{k_{2}-1}|i\rangle$, prepared exactly with no ancilla qubits and no post-selection.
An alternative construction via LCU with two STEP components is possible but requires an ancilla qubit and post-selection; the Draper-adder composition is deterministic and ancilla-free.
3.3.2 Gate count
The Draper adder dominates at $\mathcal{O}(m^{2})$ due to the QFT pair, giving $\mathcal{O}(m^{2})$ total gates in general. Two special cases admit $\mathcal{O}(m)$ circuits:
[…]
  • Power-of-2-aligned block ($w=2^{p}$, $k_{1}$ a multiple of $w$): the shift requires only $X$ gates on the upper qubits and $H^{\otimes p}$ on the lower qubits.
Example.
circuit, info = encode(SQUARE(k1=2, k2=6, c=1.0), N=8)
# info.complexity -> "O(m^2)"
# info.gate_count -> 7 (STEP(4) + Draper adder(2), m=3)
Figure 4 shows the interval vector and circuit.
Figure 4: Left: Square: $f_{[2:6]}=1$, N=8N=8Right: PyEncode circuit.
3.4 Fourier
A superposition of $T$ sinusoidal modes, each parameterized by frequency $n_{t}$, amplitude $A_{t}$, and phase $\varphi_{t}$:
$f_{i}=\sum_{t=1}^{T}A_{t}\sin\!\left(\frac{2\pi n_{t}i}{N}+\varphi_{t}\right)$ (4)
Constructor: FOURIER(modes=[(n1, A1, phi1), ...]). Circuit complexity: $\mathcal{O}(m^{2})$ via the inverse Quantum Fourier Transform [9, 20]. The DFT of each mode contributes exactly two nonzero coefficients (a complex conjugate pair at $\pm n_{t}$), so the full vector has $2T$ nonzero DFT coefficients. PyEncode prepares this sparse Fourier state using the SPARSE synthesizer and applies the inverse QFT, yielding an $\mathcal{O}(m^{2})$ circuit dominated by the QFT for $T\ll m$; the SPARSE sub-circuit contributes $\mathcal{O}(Tm)$ gates, which is subdominant when $T=\mathcal{O}(1)$. The single-mode case ($T=1$) subsumes sine ($\varphi=0$) and cosine ($\varphi=\pi/2$) as special cases of the same circuit. Arises as a sinusoidal or multi-modal load in mechanics, an eigenmode in structural dynamics, and a Fourier component in signal processing.
Example. A pure sine wave is a single Fourier mode with $\varphi=0$. The DFT has two nonzero coefficients; PyEncode prepares a sparse Fourier state and applies the inverse QFT:
circuit, info = encode(FOURIER(modes=[(1, 1.0, 0)]), N=16)
# info.complexity -> "O(m^2)"
Figure 5 shows the sinusoidal vector and inverse-QFT circuit.
Figure 5: Left: Fourier ($T=1$): $\sin(2\pi i/N)$, N=16N=16Right: PyEncode circuit.
3.5 Walsh
A two-level piecewise-constant state with period $P=2^{k+1}$:
$f_{i}=\begin{cases}c_{+}&b_{k}(i)=0\\ c_{-}&b_{k}(i)=1\end{cases}$ (5)
where $b_{k}(i)$ denotes the $k$-th bit of $i$. Constructor: WALSH(k, c_pos, c_neg). Circuit complexity: $\mathcal{O}(m)$ ($m+1$ gates).
The circuit is a single-qubit rotation on qubit $k$ followed by $H^{\otimes m}$ [27]. The rotation angle is
$\theta=2\,\mathrm{atan2}(c_{+}-c_{-},\;c_{+}+c_{-}),$ (6)
chosen so that $R_{y}(\theta)|0\rangle=\cos(\theta/2)|0\rangle+\sin(\theta/2)|1\rangle$ distributes amplitude in the ratio $c_{+}:c_{-}$ after the Hadamard layer. When $c_{-}=-c_{+}$ (the default), $\theta=\pi$ and $R_{y}(\pi)=X$, recovering the standard Walsh function (signed uniform superposition). The general case prepares a biased square wave without ancilla qubits and at the same $m+1$ gate cost. Arises in quantum signal processing, spin-chain eigenstates, and as the PREP state for two-valued Hamiltonian coefficient vectors such as the Fermi-Hubbard model — replacing a two-component LCU construction with a single ancilla-free $\mathcal{O}(m)$ circuit.
Example. The standard form ($c_{-}=-c_{+}$) is a signed uniform superposition; the generalized form encodes two distinct levels without ancilla. Both use the same $m+1$-gate circuit [27]:
# Two distinct levels, no ancilla
circuit, info = encode(WALSH(k=2, c_pos=1.0, c_neg=4.0), N=8)
# f = [1,1,1,1,4,4,4,4] / ||f||
# info.gate_count -> 4 (still m+1)
Figure 6 shows the signed superposition and $m+1$-gate circuit.
Figure 6: Left: Walsh $k=1$ (standard): period $P=4$, N=8N=8Right: PyEncode circuit.
3.6 Geometric
An exponential decay (or growth) sequence:
$f_{i}=c\,r^{\,i},\qquad 0<r,\;r\neq 1$ (7)
Constructor: GEOMETRIC(ratio, c). Circuit complexity: $\mathcal{O}(m)$ — exactly $m$ gates, zero two-qubit gates, and circuit depth 1 (all rotations execute in parallel).
3.6.1 Construction
The product-state structure of geometric sequences is a direct consequence of the binary representation of the index and has been used as an intermediate construction in prior work [28]. The vector $r^{i}$ is multiplicatively separable over the bits of $i$. Writing $i=\sum_{j=0}^{m-1}b_{j}\,2^{j}$ gives
$r^{\,i}=\prod_{j=0}^{m-1}r^{\,b_{j}\,2^{j}}.$ (8)
Each factor depends on a single bit $b_{j}$, so the normalized quantum state is a product state:
$|\psi\rangle\propto\bigotimes_{j=0}^{m-1}\bigl(|0\rangle+r^{\,2^{j}}|1\rangle\bigr).$ (9)
Each qubit $j$ is prepared independently by $R_{y}(\theta_{j})|0\rangle$ with $\theta_{j}=2\arctan(r^{\,2^{j}})$. The full circuit is $m$ independent single-qubit rotations with no entangling gates and no ancilla qubits. PyEncode surfaces this as a first-class named pattern with a validated synthesizer, making it immediately composable with the LCU framework. Arises as a discounted cash-flow model in quantum finance, as an exponentially decaying source term in diffusion problems, and as a discretized probability distribution (geometric distribution) in quantum Monte Carlo.
Example.
circuit, info = encode(GEOMETRIC(ratio=0.5), N=8)
# f = [1, 0.5, 0.25, 0.125, ...] / ||f||
# info.gate_count -> 3 (m gates, 0 CX)
# info.gate_count_2q -> 0
Figure 7 shows the exponential-decay vector and its depth-1 circuit.
Figure 7: Left: Geometric: $r=0.5$, N=8N=8Right: PyEncode circuit.
3.7 LCU (Linear Combination of Unitaries)
A weighted superposition of $r$ component states:
$|\psi\rangle\propto\sum_{j=1}^{r}w_{j}\,|\hat{f}^{(j)}\rangle,$ (10)
where each $|\hat{f}^{(j)}\rangle$ is a normalized state prepared by any PyEncode pattern. Constructor: LCU([(w1, VectorObj1), (w2, VectorObj2), …]). All weights $w_{j}>0$. Signed weights can be accommodated by absorbing the sign into the component unitary via a controlled global phase; this extension is not currently implemented.
3.7.1 Circuit construction
The circuit implements LCU [5, 3]:
  1. 1.
    PREP: apply a binary $R_{y}$-tree on $n_{\text{anc}}=\lceil\log_{2}r\rceil$ ancilla qubits, preparing $\sum_{j}\beta_{j}|j\rangle_{\text{anc}}$ with $\beta_{j}=\sqrt{w_{j}\|f^{(j)}\|}\,/\,Z$.
  2. 2.
    SELECT: apply each component circuit $U_{j}$ controlled on ancilla state $|j\rangle$.
  3. 3.
    PREP: apply the inverse $R_{y}$-tree to uncompute the ancilla.
  4. 4.
    Post-select ancilla on $|0\rangle^{\otimes n_{\text{anc}}}$.
The choice $\beta_{j}\propto\sqrt{w_{j}\|f^{(j)}\|}$ ensures that the post-selected data-register state is proportional to $\sum_{j}w_{j}|\hat{f}^{(j)}\rangle$, as required.
3.7.2 Success probability
After PREP, the amplitude in the $|0\rangle_{\text{anc}}$ subspace is $\sum_{j}\beta_{j}^{2}\,|\hat{f}^{(j)}\rangle$. The post-selection succeeds with probability:
$p=\left\|\sum_{j}\beta_{j}^{2}\,|\hat{f}^{(j)}\rangle\right\|^{2}=\sum_{i,j}\beta_{i}^{2}\,\beta_{j}^{2}\,\langle\hat{f}^{(i)}|\hat{f}^{(j)}\rangle.$ (11)
Since $\sum_{j}\beta_{j}^{2}=1$, the set $\{\beta_{j}^{2}\}$ is a probability distribution over the $r$ component states, and $p=\|\text{convex combination of unit vectors}\|^{2}\leq 1$ by the triangle inequality, with equality if and only if all component states are identical.
3.7.3 Disjoint versus overlapping support
PyEncode detects analytically whether the component vectors have disjoint support:
  • STEP and SQUARE: support is an interval; disjointness is checked by interval non-overlap in $\mathcal{O}(r^{2})$ time.
  • SPARSE: support is an index set; disjointness is checked by set intersection.
  • WALSH and FOURIER: support is always the full register $[0,N)$ — never disjoint with anything.
For disjoint-support components $\langle\hat{f}^{(i)}|\hat{f}^{(j)}\rangle=0$ for $i\neq j$, so (11) reduces to $p=\sum_{j}\beta_{j}^{4}$. For $r$ equal-weight equal-norm components, $\beta_{j}=1/\sqrt{r}$ and $p=r\cdot(1/r)^{2}=1/r$. For overlapping components, the inner products are positive, so $p$ is higher — but a UserWarning is issued since the post-selection overhead is non-trivial and amplitude amplification may be warranted.
Example. Two disjoint SQUARE intervals with different amplitudes. PyEncode detects disjoint support analytically and reports $p$ via (11):
circuit, info = encode(
LCU([(1.0, SQUARE(k1=0, k2=8, c=1)),
(3.0, SQUARE(k1=8, k2=16, c=1))]),N=16)
# f = [1,...,1, 3,...,3] / ||f||
# info.success_probability -> 0.36
For overlapping components a UserWarning is issued; post-selection is always valid and $p$ is reported in info.success_probability.
Figure 8 shows the disjoint two-interval vector and LCU circuit. Figure 9 shows an overlapping example (uniform ++ sinusoidal), where $p<1$.
Figure 8: Left: LCU (disjoint): $f_{[:8]}=1,\;f_{[8:]}=3$, N=16N=16Right: PyEncode circuit.
Figure 9: Left: LCU (overlap): uniform ++ $\sin(2\pi i/N)$, N=16N=16Right: PyEncode circuit.
3.8 Return value
encode returns (circuit, info), where circuit is a Qiskit QuantumCircuit and info is an EncodingInfo dataclass with the following fields:
  • vector_type — pattern name (e.g. "SPARSE", "GEOMETRIC").
  • N, m — vector length and number of qubits.
  • params — supplied vector parameters (e.g. {"ratio": 0.95, "c": 1.0}).
  • gate_count — total gates in the returned circuit (pre-transpilation).
  • gate_count_1q, gate_count_2qUU and CX gate counts after transpilation to $\{\textsc{cx},U\}$.
  • circuit_depth — circuit depth after transpilation; determines the minimum execution time when gates on disjoint qubits run in parallel.
  • complexity — asymptotic gate complexity (e.g. "O(m)" or "O(mˆ2)").
  • success_probability — always 1.0 for single-pattern constructors; $p\in(0,1]$ for LCU.
  • circuit_code — a human-readable Qiskit snippet that reproduces the circuit independently of PyEncode.
  • validatedTrue if statevector validation was performed (see Section 3.9).
  • vector — the classically constructed amplitude vector $\mathbf{f}$, populated only when validate=True; requires $\mathcal{O}(2^{m})$ memory.
3.9 Validation
No classical vector is constructed during synthesis — there is no vector to validate. The optional statevector check (validate=True) constructs $\mathbf{f}$ from the supplied parameters, runs the circuit on Qiskit’s statevector simulator, and verifies $\|\hat{\mathbf{f}}-\hat{\mathbf{f}}_{\mathrm{sim}}\|_{2}<\varepsilon$. This is the only validation path available and is disabled by default due to its $\mathcal{O}(2^{m})$ memory cost. For large $m$, partial verification via measurement sampling or a SWAP-test overlap estimate can provide confidence at polynomial cost; these are not currently implemented. When enabled, the constructed vector is also returned as info.vector for inspection and debugging.
circuit, info = encode(
FOURIER(modes=[(1, 1.0, 0)]), N=16,
validate=True, tol=1e-6)
# info.validated-> True
# info.vector -> numpy array of length N
4 Patterns Not Covered and Why
PyEncode produces exact, closed-form circuits for vectors that are either piecewise constant with $\mathcal{O}(1)$ pieces, sparse in the Fourier basis, or multiplicatively separable over the bits of the index (the GEOMETRIC pattern). Four practically important vector types fall outside this library. For each, approximate encoding via MPS truncation [14] or piecewise-QSVT [22] provides the natural alternative.
4.1 Ramp (linear)
The vector $f_{i}=i/N$ arises as a discretized pressure gradient or linearly tapered structural load. Although the computational basis already encodes the index $i$, routing a register value into an amplitude requires an oracle whose cost is $\mathcal{O}(2^{m})$ in the worst case, and the DFT of the ramp has $N$ nonzero coefficients, ruling out a QFT shortcut.
4.2 Poiseuille (parabolic)
The laminar pipe-flow profile $f_{i}\propto 1-(2i/N-1)^{2}$ combines a ramp term and a quadratic term, both of which resist exact compact encoding. No $\mathcal{O}(\mathrm{poly}(m))$ exact circuit is known for general parabolic profiles.
4.3 Hat (tent) function
The piecewise-linear FEM basis function centered at $i_{0}$ with half-width $h$ [25] has linearly graded amplitudes within its support, inheriting the ramp obstruction on each half. No exact $\mathcal{O}(\mathrm{poly}(m))$ circuit is known.
4.4 Gaussian
The vector $f_{i}=\exp(-\alpha(i-i_{0})^{2}/N^{2})$ appears as an initial condition in turbulence simulations and as a probability distribution in quantum Monte Carlo [19]. Its DFT is also Gaussian-shaped with $\mathcal{O}(N)$ effectively nonzero coefficients, precluding both exact encoding and a QFT shortcut. In contrast, the simple exponential $e^{-\alpha i}$ is multiplicatively separable over the bits of $i$ and admits an exact $\mathcal{O}(m)$ product-state circuit (Section 3.6); Xie and Ben-Ami [28] exploit this product state as an intermediate step for approximate Gaussian preparation via a subsequent QFT. The quadratic exponent in the Gaussian destroys this separability. Holmes and Matsuura [14] showed that smooth functions including the Gaussian admit approximate $\mathcal{O}(\mathrm{poly}(m))$ circuits via MPS truncation; this is the recommended path.
5 Gate Count Comparison
Table 2 summarizes gate counts and circuit depth at $N=64$ ($m=6$ qubits). All circuits are transpiled to $\{\textsc{cx},U\}$ (optimization_level=3, Qiskit 2.3.1). Qiskit StatePreparation uses reps=3. Two-qubit (CX) gates are hardware-critical as they are significantly noisier than single-qubit gates on near-term devices; circuit depth determines the minimum execution time when gates on disjoint qubits run in parallel. The $\mathcal{O}(m)$ patterns (Sparse, Step, Walsh, Geometric) achieve depth 1 at $m=6$, meaning all gates execute in a single parallel layer. Figure 10 shows gate count as a function of $m$. The $\mathcal{O}(m)$ patterns outperform Qiskit at all $m$. FOURIER ($\mathcal{O}(m^{2})$) crosses over at $m=5$, after which its asymptotic advantage over Qiskit’s $\mathcal{O}(2^{m})$ is clear. Gate count is independent of the number of Fourier modes $T$ (see Table 2, rows for $T=1$ and $T=2$).
Table 2: Gate counts and circuit depth at $N=64$ ($m=6$) unless noted. All circuits transpiled to $\{\textsc{cx},U\}$ (optimization_level=3, Qiskit 2.3.1). Qiskit uses general-purpose $\mathcal{O}(2^{m})$ state preparation [23]; low gate counts for Walsh and Geometric at $m=6$ reflect optimizer-level structure detection that breaks down at larger $m$ (see Geometric $m=10$ row).
PyEncode Qiskit
Pattern $U$ CX Depth $\mathcal{O}(\cdot)$ $U$ CX Depth
Sparse (s=1s\!=\!1, k=19k\!=\!19) 3 0 1 $\mathcal{O}(sm)$ 50 47 93
Sparse (s=2s\!=\!2) 4 2 3 $\mathcal{O}(sm)$ 23 19 37
Step ($k_{s}\!=\!4$) 2 0 1 $\mathcal{O}(m)$ 32 30 60
Square ($[12,52)$, general) 41 22 37 $\mathcal{O}(m^{2})$ 47 41 82
Walsh (k=1k\!=\!1, P=4P\!=\!4) 6 0 1 $\mathcal{O}(m)$ 7 1 3
Geometric (r=0.95r\!=\!0.95, m=6m\!=\!6) 6 0 1 $\mathcal{O}(m)$ 6 0 1
Geometric (r=0.95r\!=\!0.95, m=10m\!=\!10) 9 0 1 $\mathcal{O}(m)$ 1,022 1,013 2,027
Fourier (T=1T\!=\!1, n=1n\!=\!1, $\varphi\!=\!0$) 45 33 41 $\mathcal{O}(m^{2})$ 63 57 115
Fourier (T=1T\!=\!1, n=3n\!=\!3, $\varphi\!=\!\pi/4$) 45 35 44 $\mathcal{O}(m^{2})$ 63 57 115
Fourier (T=2T\!=\!2) 48 35 43 $\mathcal{O}(m^{2})$ 60 57 112
Figure 10: Transpiled gate count vs. number of qubits $m$ ($N=2^{m}$), for $m=4,6,8,10,12$. All PyEncode patterns remain below 30 gates across the full range, while Qiskit StatePreparation on a random vector grows from 26 gates at $m=4$ to 8,178 at $m=12$. SPARSE uses $s=2$ entries with non-aligned indices; SQUARE uses a non-aligned interval [N/4+1, 3N/4+1)[N/4{+}1,\;3N/4{+}1), which activates the general $\mathcal{O}(m^{2})$ Draper-adder path. All patterns scale as $\mathcal{O}(\mathrm{poly}(m))$. Circuit depth follows the same asymptotic scaling; see Table 2 for per-pattern depth at $m=6$.
All circuits in this work are exact in the noiseless setting. Characterizing fidelity under realistic noise models (e.g., depolarizing or device-calibrated noise) is an important direction for future work; the depth-1 circuits (Walsh, Geometric, single-entry Sparse, Step) are expected to be particularly robust since they contain no two-qubit gates.
6 Applications
This section demonstrates the PyEncode framework through three applications drawn from distinct fields.
6.1 Quantum Chemistry
Fault-tolerant quantum algorithms for chemistry represent the molecular Hamiltonian as a linear combination of unitaries (LCU), $H=\sum_{j}\alpha_{j}\hat{P}_{j}$ [4]. This requires a PREP oracle that prepares $|\boldsymbol{\alpha}\rangle\propto\sum_{j}\sqrt{|\alpha_{j}|}\,|j\rangle$ [5], whose gate cost directly affects the total $T$-gate count [3].
Fermi-Hubbard model. After a Jordan-Wigner transformation on an $L$-site chain [15, 16], the Pauli coefficient vector takes exactly two distinct values: the hopping amplitude $t$ on the first $L$ terms and the on-site interaction $U$ on the remaining $L$ terms ($N=2L$ total). This two-valued equal-block structure is exactly a generalized Walsh function, giving an ancilla-free single-register PREP circuit in $m+1=\log_{2}(2L)+1$ gates:
import math
L = 8; t = 1.0; U = 4.0
circuit, info = encode(
WALSH(k=int(math.log2(L)), c_pos=t, c_neg=U),
N=2*L)
# info.complexity -> "O(m)"
# info.gate_count -> 5 (log2(16)+1)
Figure 11 shows the coefficient vector and circuit. Although the coefficient vector resembles two concatenated step functions, the generalized Walsh synthesizer prepares it in a single $m+1$-gate circuit without ancilla or post-selection.
Figure 11: Left: Fermi-Hubbard PREP: t=1t\!=\!1, U=4U\!=\!4, L=8L\!=\!8, N=2L=16N\!=\!2L\!=\!16Right: PyEncode circuit.
This yields an $\mathcal{O}(m)$ PREP circuit with $m=\log_{2}(2L)$ qubits and no ancilla — a direct consequence of the equal-block two-valued structure that the generalized Walsh synthesizer exploits.
Heisenberg spin chain. For translationally invariant nearest-neighbour Hamiltonians ($XXZ$, Ising), all exchange coefficients are equal — a uniform coefficient vector, which is the full-range special case STEP(k_s=N):
circuit, info = encode(STEP(k_s=L, c=J), N=L)
In both cases, the gate saving over Qiskit’s general preparation is exponential, and it applies to every PREP–SELECT repetition required by phase estimation [3].
In passing, we note that the Pauli coefficient vector of the 1D Poisson stiffness matrix also admits efficient encoding: its magnitudes follow a geometric series $2^{-k}$, directly handled by the GEOMETRIC pattern.
6.2 Computational Mechanics
The Poisson equation $\nabla^{2}u=f$ on the unit square with a separable sinusoidal source,
$f(x,y)=\sin(2\pi nx)\sin(2\pi py),$ (12)
arises naturally in elliptic PDE solvers when the source term is periodic [25]. After discretization on an $N\times N$ grid, the right-hand side vector $\mathbf{f}$ is a tensor product:
$\mathbf{f}=\mathbf{u}\otimes\mathbf{v},u_{i}=\sin(2\pi ni/N),v_{j}=\sin(2\pi pj/N).$ (13)
A tensor product of two normalized states is a product state on the qubit register, so the $2m$-qubit encoding separates exactly into two independent $m$-qubit FOURIER circuits:
circ_x, _ = encode(FOURIER(modes=[(2, 1.0, 0)]), N=32)
circ_y, _ = encode(FOURIER(modes=[(3, 1.0, 0)]), N=32)
circuit = circ_x.tensor(circ_y)
Figure 12 shows the separable source term and combined circuit.
Figure 12: Left: 2D Poisson source: $\sin(4\pi i/N)\sin(6\pi j/N)$, N=32N=32Right: PyEncode circuit.
The total gate count is $2\times\mathcal{O}(m^{2})$ for $2m$ qubits encoding $N^{2}$ amplitudes, compared to $\mathcal{O}(N^{2})$ for general state preparation.
6.3 Quantum Finance
Quantum amplitude estimation (QAE) offers a quadratic speedup over classical Monte Carlo for computing expectations $\mathbb{E}[f(X)]$ [19]. The state preparation oracle must load the discretized probability distribution into a quantum register. A piecewise-constant approximation — where the price axis is divided into bins of equal width, each with a distinct probability mass — is a natural fit for the LCU constructor with disjoint SQUARE components:
circuit, info = encode(
LCU([(0.10, SQUARE(k1=0, k2=4, c=1.0)),
(0.40, SQUARE(k1=4, k2=8, c=1.0)),
(0.35, SQUARE(k1=8, k2=12, c=1.0)),
(0.15, SQUARE(k1=12, k2=16, c=1.0))]),
N=16)
# info.success_probability -> 0.315
PyEncode detects that the four intervals are disjoint and computes $p=0.315$ analytically via (11). Figure 13 shows the 4-bin distribution and the LCU circuit.
Figure 13: Left: Price distribution as LCU of 4 disjoint SQUARE bins, N=16N=16Right: PyEncode circuit.
Since QAE applies the state preparation oracle $\mathcal{O}(1/\varepsilon)$ times [19], the $\mathcal{O}(\mathrm{poly}(m))$ circuit cost of each LCU component compounds into substantial savings over a general $\mathcal{O}(2^{m})$ preparation.
7 Conclusions
PyEncode provides an open-source implementation of theoretical work on structured quantum state preparation that until now lacked runnable code. The library exposes a single function encode(VectorObj, N) that maps a typed parameter declaration directly to a closed-form Qiskit circuit, without code parsing and without materializing a classical vector.
The library implements the following pattern synthesizers, each mapping a typed parameter declaration to an exact, closed-form Qiskit circuit with no vector materialization:
  • SPARSE — Gleinig–Hoefler $\mathcal{O}(sm)$ sparse state preparation [8].
  • STEP — Shukla–Vedula $\mathcal{O}(m)$ prefix uniform superposition [24].
  • SQUARE — general interval $k_{1},k_{2})$ via STEP(w)(w) composed with a Draper QFT-based constant adder [6]; $\mathcal{O}(m^{2})$ in general, $\mathcal{O}(m)$ for power-of-2-aligned blocks.
  • WALSH — two-level piecewise-constant state via $R_{y}(\theta)+H^{\otimes m}$ [27], with the generalized form supporting asymmetric levels $c_{+}\neq-c_{-}$.
  • GEOMETRIC — exponential decay / geometric sequence $f_{i}=c\,r^{\,i}$ as a product state [28]; $\mathcal{O}(m)$ gates, zero two-qubit gates.
  • FOURIER — exact sinusoidal state preparation via inverse QFT [9, 20], at $\mathcal{O}(m^{2})$ cost independent of the number of modes.
  • LCU — LCU [3] for weighted superpositions of $r$ component states, with analytically computed success probability and automatic disjoint-support detection.
Several vector types common in practice — ramp, hat, Gaussian — do not admit exact $\mathcal{O}(\mathrm{poly}(m))$ circuits and fall outside the pattern library; approximate encoding via MPS truncation [14] or piecewise-QSVT [22] is recommended for these cases, and is planned for a future pyencode.approx submodule. In particular, the Walsh Series Loader of Zylberman and Debbasch [29] provides a natural foundation for approximate encoding of smooth functions via truncated Walsh–Hadamard series, extending the single-term exact construction already implemented in WALSH.
Characterizing which classes of vectors admit $\mathcal{O}(\mathrm{poly}(m))$ exact circuits remains an open and potentially rich theoretical question. The pattern library presented here — six closed-form families and one composition rule — provides a concrete starting point.
Acknowledgments
The first author would like to acknowledge the Vilas Associate Grant from the University of Wisconsin Graduate School.
Use of AI Tools
The authors used generative AI assistants including Claude (Anthropic) and Gemini (Google) during the preparation of this manuscript. The tools were used for code development and drafting of text. All content was reviewed, verified, and edited by the authors, who take full responsibility for the accuracy and integrity of the work.
Declarations
The authors declare no conflict of interest.
Code Availability
The code developed in this work is available at https://github.com/UW-ERSL/PyEncode.git.